An object of mass 0.5 kg is whirled at the end of string 0.8 m long. If the string makes three revolutions in 1.2 sec, find the tension in the string.

Grade:11

2 Answers

Nirmal Singh.
9 years ago
when object will move whirkled it will move in circular move. to move in circular motion tension will provide the centripetal force.

So T = mv2/r

and v = 2pi r / t

where t is the time period in one revolution t = 1.2 / 3 = 0.4 sec

T = m *(2pi r / t)2/ r

T = 4pi2r m / t2

T = (4*3.14*3.14 *0.8*0.5 ) / (0.4*0.4)

T = 98.596 N

Regards,
Nirmal Singh
Askiitians Faculty
Pawan Kumar
5 years ago
An object of mass 0.5 kg whirled at the end of string o.8 m long. If the string makes revolution in 1.2 . Find tension in the string